OpenStudy (anonymous):

mean = 75; variance = 18.49, if a person is selected at random, what is the probability that their resting pulse rate is under 70?

OpenStudy (anonymous):

Assuming a normal distribution:\[\sigma=\sqrt{18.49}=4.3\]\[Z=\frac{70-75}{4.3}\approx-1.16\]\[P(Z<-1.16)=1-P(Z<1.16)=1-\Phi(1.16)=1-0.877=0.123\]

OpenStudy (anonymous):

why can't you just use normal cdf (0,70,75,18.490?

OpenStudy (anonymous):

You can, that's what I did. Maybe you're not familiar with the notation?

OpenStudy (anonymous):

oh you know what .. i assumed that 18.49 was the standard deviation but you square rooted it... why do you square root it for? is it because it is the "variance?"

OpenStudy (anonymous):

Yes, variance is defined as standard deviation squared.

OpenStudy (anonymous):

ok thanks for clarifying that for me :) i get it now

OpenStudy (anonymous):

Glad to hear it.